### v1

# Butter: Frequency Consistency and Hierarchical Fusion for Autonomous Driving Object Detection

Abstract.
Hierarchical feature representations play a pivotal role in computer vision, particularly in object detection for autonomous driving. Multi-level semantic understanding is crucial for accurately identifying pedestrians, vehicles, and traffic signs in dynamic environments. However, existing architectures, such as YOLO and DETR, struggle to maintain feature consistency across different scales while balancing detection precision and computational efficiency. To address these challenges, we propose Butter, a novel object detection framework designed to enhance hierarchical feature representations for improving detection robustness. Specifically, Butter introduces two key innovations: Frequency-Adaptive Feature Consistency Enhancement (FAFCE) Component, which refines multi-scale feature consistency by leveraging adaptive frequency filtering to enhance structural and boundary precision, and Progressive Hierarchical Feature Fusion Network (PHFFNet) Module, which progressively integrates multi-level features to mitigate semantic gaps and strengthen hierarchical feature learning. Through extensive experiments on BDD100K, KITTI, and Cityscapes, Butter demonstrates superior feature representation capabilities, leading to notable improvements in detection accuracy while reducing model complexity. By focusing on hierarchical feature refinement and integration, Butter provides an advanced approach to object detection that achieves a balance between accuracy, deployability, and computational efficiency in real-time autonomous driving scenarios. Our model and implementation are publicly available at https://github.com/Aveiro-Lin/Butter, facilitating further research and validation within the autonomous driving community.
Hierarchical feature representations; Multi-scale feature fusion; Object detection; Autonomous driving; Frequency-adaptive feature enhancement
conference: the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandccs: Computing methodologies Hierarchical representations are equally corresponding authors.
The paper is supported by National Natural Science Foundation of China No.62472046.
Figure 1. Comparison between the neck of proposed Butter and other 2 previous most popular 2D object detection methods YOLOv12 (Tian et al., 2025) and Hyper-YOLO (Feng et al., 2024).
1. Introduction
Recent advances in autonomous driving technology have made object detection essential for intelligent transportation systems. This technology allows for real-time detection of objects such as pedestrians, cyclists, and traffic signs, which is crucial for decision-making (Mahaur and Mishra, 2023; Wang et al., 2025b; Benjumea et al., 2021; Wang et al., 2025a; Zhang et al., 2025c). Despite these advancements, challenges exist due to the complexity of traffic scenarios and diverse targets. The need for real-time performance exposes limitations in the accuracy, robustness, and real-time capabilities of current algorithms in dynamically changing environments (Chen et al., 2021; Wang et al., 2023b).
Object detection algorithms in deep learning are typically categorized into two types: two-stage and one-stage algorithms. One-stage detectors, such as YOLO (You Only Look Once) models (Glenn, 2023; Wang et al., 2024a; Khanam and Hussain, 2024; Tian et al., 2025; Feng et al., 2024; Wang et al., 2023a, 2024b) and SSD (Liu et al., 2016), are known for their high-speed detection capabilities. In contrast, two-stage detectors like R-CNN (Girshick et al., 2016) and faster R-CNN (Ren et al., 2017) tend to offer higher detection accuracy. Although two-stage models excel in accuracy, one-stage models are preferred for their faster processing times. In response to the real-time demands of autonomous driving scenarios, we focus lightweight approaches, which fall into the category of one-stage object detection.
Current novel object detection algorithms, such as Hyper-YOLO (Feng et al., 2024) and YOLOv12 (Tian et al., 2025), have demonstrated strong performance in detecting large objects in autonomous driving. However, they still face challenges when it comes to handling complex backgrounds, occlusions, and objects with varying scales (Zamanakos et al., 2021; Ding et al., 2024). This is especially evident in traffic environments, where factors like differences in object sizes, changes in lighting, and background noise further complicate detection tasks (Gupta et al., 2021; Huang et al., 2025; Zhang et al., 2025a). Moreover, effectively detecting objects in autonomous driving systems, which require high real-time performance, remains a significant challenge for existing technologies (Bachute and Subhedar, 2021; Tilijani et al., 2023). To address this, object detection algorithms must not only excel at precise feature extraction, but also be optimized for computational efficiency and designed to be lightweight, ensuring that they meet the real-time and high-efficiency demands of autonomous driving systems while being easy to deploy (Liu et al., 2021; Subramanian et al., 2023).
To achieve a balance between lightweight design and detection accuracy, we propose a novel 2D real-time object detection model, Butter. This model achieves real-time object detection with high accuracy while maintaining a low parameter count for efficient deployment. It reduces computational overhead, making it ideal for autonomous driving scenarios. As shown in Fig. 1, unlike the latest advanced models such as Hyper-YOLO (Feng et al., 2024) and YOLOv12 (Tian et al., 2025), our model incorporates two modules in the neck: the Frequency-Adaptive Feature Consistency Enhancement (FAFCE) and the Context-Aware Spatial Fusion (CASF). These modules optimize feature alignment and reduce semantic gaps between feature levels, thereby enhancing category consistency and boundary precision while improving the multi-scale feature representation capability of the object detection model. In parallel, recent works have explored using language priors to enhance feature expressiveness and open-set robustness in road anomaly detection (Tian et al., 2023). In experiments, our lightweight Butter method with low parameters achieves accurate detection on 3 autonomous driving datasets.
Our main contributions are summarized as follows.
  • We propose a FAFCENet component, which employs contextual low-frequency damping to refine semantic representations, contextual high-frequency amplifier to restore object boundaries, and a feature resampling module to improve spatial consistency, thereby enhancing category coherence and boundary precision for robust object localization and classification.
  • We propose PHFFNet, which progressively integrates hierarchical feature aggregation to enhance semantic representations. It employs CASF to optimize feature alignment, reducing semantic gaps across hierarchical levels and enhancing multi-scale feature representation in object detection.
  • Butter addresses the limitations of existing object detection methods, particularly in integrating multi-scale features within the neck layer. The model achieves state-of-the-art (SOTA) results with fewer than 10 million parameters on datasets like KITTI (Geiger et al., 2012), Cityscapes (Cordts et al., 2016), and BDD100K (Yu et al., 2020). For instance, on the Cityscapes (Cordts et al., 2016) dataset, Butter reduces the parameter count by over 40% while improving mAP@0.5 by 1.8%, showcasing significant advances in both detection accuracy and parameter-efficiency.
2. Related Works
2.1. Frequency-Based Feature Refinement.
Frequency analysis, a fundamental technique in traditional signal processing (Pitas, 2000), has proven valuable in deep learning, especially for computer vision. It has been used to investigate generalization abilities (Wang et al., 2020) and optimization techniques (Yin et al., 2019) in Deep Neural Networks (DNNs). Manipulating high-frequency components enables adversarial attacks (Luo et al., 2022; Yin et al., 2019) and weaken feature representation by reducing intra-category similarity (Luo et al., 2022). Studies by Rahaman et al. (Rahaman et al., 2019) and Xu et al. (Xu and Zhou, 2021) show that DNNs prioritize low-frequency patterns during training, a phenomenon known as spectral bias. Zhang et al. (Zhang, 2019) examine the effects of frequency aliasing on translation invariance, while FLC (Grabinski et al., 2022) confirms that aliasing reduces model robustness. Low-pass filters, as demonstrated by Chen et al. (Chen et al., 2023), can suppress high-frequency noise, particularly in low-light conditions, and AdaBlur (Zou et al., 2023) mitigates aliasing by applying content-aware low-pass filtering during downsampling. Further optimization, leveraging additional frequency components, has shown to improve performance (Qin et al., 2021; Magid et al., 2021), particularly through discrete cosine transform coefficients for channel attention mechanisms. Traditional convolutional theorems in DNNs (Huang et al., 2023) use adaptive frequency filters to act as global label mixers, and various frequency-domain methods have been integrated into DNNs to enhance non-local feature learning (Huang et al., 2023; Rao et al., 2021; Li et al., 2021).
Current frequency analysis improves neural network generalization, adversarial robustness, and feature extraction but focuses on single-frequency operations, neglecting frequency consistency in multiscale fusion. In contrast, the FAFCE component integrates low-frequency damping, high-frequency amplifier, and displacement calculator to ensure frequency consistency across levels, adapting to data variations, reducing information loss, and enhancing multiscale detection and localization.
2.2. Feature Fusion.
Feature fusion enhances both semantic and spatial representations by integrating low- and high-resolution features. This is achieved through top-down approaches such as DeepLabv3+ (Chen et al., 2018) and U-Net (Falk et al., 2019), as well as bottom-up approaches like SeENe (Pang et al., 2019) and DLA (Yu et al., 2018). However, simple fusion methods struggle with resolution and semantic inconsistencies. Modern techniques address these challenges through sampling-based methods, such as SFNet (Li et al., 2023b), FaPN (Huang et al., 2021a), and AlignSeg (Huang et al., 2021b), which align features using spatial offsets. They also include kernel-based methods, such as deconvolution (Zeiler and Fergus, 2014), Pixel Shuffle (Shi et al., 2016), A2U (Dai et al., 2021), and CARAFE (Wang et al., 2021), which upscale features using fixed or learnable kernels. Fusion modules like CARAFE (Wang et al., 2021), ASFF (Liu et al., 2019), and DRFPN (Ma and Chen, 2020) improve multi-scale feature fusion within FPN, enhancing detection accuracy (Jin et al., 2024).
FPN-based methods (Lin et al., 2017) improve object detection by predicting multi-level features, but they struggle to integrate high- and low-level features effectively. To address this, PANet (Liu et al., 2018) adds a bottom-up pathway, and NASFPN (Ghiasi et al., 2019) uses neural architecture search to optimize feature connections. Zhao et al. (Zhao et al., 2019), Wu et al. (Wu et al., 2023), and Ma et al. (Ma et al., 2023) propose alternative optimization methods. While GraphFPN (Zhao et al., 2021) and FPT (Zhang et al., 2020) employ graph neural networks and self-attention (Li et al., 2022) for feature exchange and aggregation, respectively, they increase computational complexity. In contrast, PHFFNet uses standard convolutions for a more efficient solution, making it more suitable for real-world applications. Similar to multimodal transformers in embodied AI (Li et al., 2023a), our design emphasizes structured feature alignment with minimal overhead.
Figure 2. Comprehensive workflow of the Butter model for autonomous driving object detection. (1) The workflow begins with a monocular image of $640\times 640$ pixels processed through the HGStem of the Backbone to extract features. These features are subsequently refined through a series of lightweight HGBlocks, Depthwise Convolutions (DWConv) (Howard et al., 2017), and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) before entering the Neck module. The Neck consists of two parts: FAFCENet and PHFFNet. Following the Neck, the model utilizes four heads in the Head layer to produce outputs, including the original image annotated with class labels, confidence scores, and bounding boxes. (2) The CBAM module in the lower left corner applies channel and spatial attention to guide the network toward informative features. (3) The Hierarchical Fusion Block in the upper right corner enables multi-level feature interaction within the Context-Aware Spatial Fusion (CASF) module. Horizontal arrows represent feature exchange, while only diagonal arrows are used to indicate upsampling and downsampling.
Current feature fusion methods improve high-low feature integration but struggle with cross-level inconsistency, boundary loss, and high computational cost. PHFFNet uses hierarchical aggregation and CASF for efficient integration. The FAFCE component integrates adaptive damping, high-frequency amplification, and feature resampling to improve spatial and frequency consistency, thereby enhancing feature representation and boosting multi-scale detection performance.
3. Method
3.1. Overview
Task Description. Object detection in autonomous driving is crucial for identifying various objects, such as pedestrians, vehicles, and traffic signs, from monocular RGB images in complex, dynamic environments. We present Butter, a novel object detection method specifically designed for autonomous driving scenarios.
Overall Framework. As shown in Fig.2, Butter consists of three primary branches. The Backbone Branch (Fig.2 (a)) leverages a lightweight version of the HGNetV2 architecture, enhanced with lightweight Depthwise Convolutions (DWConv) (Howard et al., 2017) and Convolutional Block Attention Modules (CBAM) (Woo et al., 2018) Module. This branch processes input monocular $640\times 640$ RGB images to extract essential features, improving both computational efficiency and detection accuracy. The Neck Branch (Fig. 2 (b)) consists of two crucial elements: FAFCE as a component and PHFFNet as a module. The FAFCE component enhances the model’s ability to maintain feature consistency and boundary precision. At the same time, the PHFFNet module integrates low-level and high-level features progressively, helping to bridge semantic gaps and improve the accuracy of object detection. The Head Branch (Fig. 2 (c)) distinguishes itself from traditional YOLOv12 (Tian et al., 2025) models by using four detection heads instead of the typical three.
3.2. Butter Method
Backbone Branch. HGNetV2 is short for PP-HGNetV2, a new network model developed by Baidu PaddlePaddle Vision Team. Due to its excellent real-time performance and accuracy, it has shown outstanding results in tasks such as single- or multilabel classification, object detection, and semantic segmentation. We use it as the baseline for our backbone model.
Figure 3. Architecture comparison between the original HGNetV2 and the lightweight backbone of Butter.
We make lightweight improvements based on the original HGNetV2, as illustrated in Fig. 3, the first step of Butter involves processing monocular RGB images through an HGStem to extract features. These features are then passed through the first stage of the lightweight HGBlock. The main difference between the lightweight HGBlock we propose and the traditional HGBlock is that we replace the convolutional layers with lightweight layers such as GhostConv, RepConv, DWConv, and LightConv, thereby reducing the parameter count of the model’s backbone. Compared to the traditional HGNetV2, we have made the following optimizations. We replace the LDS (Learnable Down-sampling) operation that was originally applied starting from Stage 2-4 with DWConv. The DWConv has been widely applied in various object detection models. Additionally, after Stage 4, we replace the Global Average Pooling (GAP) with Spatial Pyramid Pooling Fast (SPPF) and substitute the Fully Connected (FC) layer with CBAM. These adjustments improve the model’s computational efficiency and enhance its feature extraction and attention mechanisms. The introduction of SPPF and CBAM into the backbone, rather than the neck or head, leverages the role of the backbone in the early extraction of features. In autonomous driving, optimizing feature quality and multi-scale information at this stage is critical for perceiving complex environments. By placing these modules here, we enhance the features’ discriminative power, improving downstream tasks like object localization and classification. This strategy boosts overall network performance, ensuring robust feature extraction for accurate and efficient real-time detection in driving scenarios.
The decision to add SPPF and CBAM after Stage 4, rather than earlier, stems from the distinct roles of each stage in object detection. Stages 1 to 3 focus on extracting basic and fine-grained features, emphasizing low-level structures from the input image, without considering object scales or specific feature selection. In autonomous driving, this early extraction is critical for fast, accurate detection of objects in dynamic environments. Introducing SPPF and CBAM at this stage would add unnecessary computational load and disrupt the learning of low-level features. Stage 4, being the final stage, contains rich contextual information and multi-level features, which are crucial for precise object classification and localization. Therefore, placing SPPF and CBAM after Stage 4 enables the refinement and fusion of multi-scale features, enhancing the network’s ability to handle complex driving scenarios with higher precision.
Neck Branch. The Neck module is composed of two key components. After the feature map passes through the backbone, it enters the FAFCENet, which consists of multiple FAFCE components and convolutional (Conv) layers. The FAFCE module is designed to improve feature fusion by ensuring the consistency and accuracy of multi-level features, while the convolutional layers further refine the features. By stacking multiple FAFCE components and applying convolutional layers, FAFCENet progressively refines feature information, enabling effective fusion of features across different levels. This fusion is crucial for handling the complex, multi-scale nature of autonomous driving environments. The PHFFNet then combines low-level and high-level features through the Hierarchical Fusion Block, addressing semantic discrepancies between non-adjacent layers and enhancing feature integration. The CASF mechanism in the blocks resolves conflicts between layers, ensuring the retention of critical information for accurate object detection. Finally, the refined feature map is passed into the Head layer for classification and localization. For details of the FAFCE and PHFFNet, refer to Sec. 3.3 and 3.4.
Head Branch. The Head module is responsible for receiving the feature map from the Neck and performing tasks such as object classification and bounding box regression. The Butter model uses four heads instead of the traditional three or more heads to balance multi-task processing and computational efficiency. The four heads allow for more detailed handling of multiple tasks in complex scenarios (such as object detection, lane line recognition, etc.), while avoiding the computational burden that comes with five or more heads. On the other hand, three heads may not fully meet the diverse demands of autonomous driving environments. Therefore, the use of four heads improves detection accuracy while maintaining good efficiency.
3.3. Frequency-Adaptive Feature Consistency Enhancement (FAFCE) Component
Current object detection models face challenges in hierarchical representation learning, mainly due to the lack of semantic information in low-level features and spatial loss in high-level features. Traditional fusion methods often cause information loss and inconsistencies, impairing boundary detection and small object recognition, which are vital in autonomous driving. Inspired by how shape priors enhance zero-shot segmentation (Liu et al., 2023), the FAFCE component employs frequency-aware damping and amplification to optimize the fusion of low-level and high-level features. This improves multi-scale information capture, enhancing detection accuracy and robustness in the dynamic, complex environments of autonomous driving.
We present the details of the FAFCE component, as illustrated in Fig. 7 in the Supplementary Material, where the figure is included due to space constraints. FAFCE operates in 3 main stages: preliminary fusion, resampling and refined fusion. Before delving into the three stages, we will first provide an overview of traditional feature fusion methods in image processing, focusing on their formulations, which will set the stage for a detailed comparison with our proposed method.
A widely adopted approach to feature fusion can be expressed as:
(1) $\mathbf{B}^{l}=\mathbf{A}^{l}+\mathcal{U}_{\theta}^{\mathrm{UP}}(\mathbf{B}^{l+1}),$
where $\mathbf{A}^{l}\in\mathbb{R}^{2H\times 2W\times C}$ represents the $l$-th feature map produced by the backbone, and $\mathbf{B}^{l+1}\in\mathbb{R}^{H\times W\times C}$ denotes the fused feature at the $(l+1)$-th level. The operator $\mathcal{U}_{\theta}^{\mathrm{UP}}$ refers to an upsampling function and $\theta$ is the learnable parameter. The two feature maps are assumed to have the same number of channels; otherwise, a $1\times 1$ convolution can be applied as a projection function, which we omit here for simplicity. Despite its simplicity, this fusion approach causes semantic inconsistency and boundary misalignment. These issues harm dense prediction in autonomous driving. Simple interpolation spreads errors spatially, which leads to misclassifications. It also oversmooths outputs, weakening boundary localization. Additionally, the approach underutilizes boundary cues from lower-level features, which are vital for accurate detection in complex environments.
As shown in Figure 7, the proposed FAFCE consists of 2 important modules: the high-frequency amplifier, and the low-frequency damping. The 2 modules can be written as follows:
High-Frequency Amplifier. The amplifier amplifies high-frequency components of the input features, improving the fine details of the object boundaries. It is formulated as:
(2) $\widetilde{\mathbf{A}}^{l}=\mathbf{A}^{l}+\mathcal{H}^{\mathrm{HF}}_{\phi}\left(W^{\mathrm{HF}}_{\mathrm{l}}\odot\mathbf{A}^{l}\right),$
where $\mathcal{H}^{\mathrm{HF}}_{\phi}$ represents the high-frequency amplifier operation with learnable parameters $\phi$, $W^{\mathrm{HF}}_{\mathrm{l}}$ is the learnable filter matrix for high-frequency amplifier, and $\odot$ denotes element-wise multiplication, enhancing the high-frequency features to provide better resolution for fine boundaries.
Low-Frequency Damping: The damping operation suppresses low-frequency components that may introduce noise into the feature map. The operation is formulated is
(3) $\widetilde{\mathbf{B}}^{l+1}=\mathcal{U}_{\theta}^{\mathrm{UP}}\left(\mathcal{L}_{\psi}^{\mathrm{LF}}\left(\mathbf{B}^{l+1}\right)\right).$
The $\mathcal{L}_{\psi}^{\mathrm{LF}}$ represents the low-frequency damping operation with learnable parameters $\psi$, which helps suppress irrelevant low-frequency features. $\mathcal{U}_{\theta}^{\mathrm{UP}}$ is the upsampling operator, restoring the spatial resolution of the feature map after damping. This operation ensures that the model focuses on the high-frequency details and ignores the irrelevant low-frequency components, leading to sharper and more accurate predictions.
Three Stages of FAFCE. These stages work sequentially to refine and enhance the feature maps in FAFCE, ultimately leading to the final fused features used for prediction.
Stage 1: Preliminary Fusion
In Stage 1, the initial fusion of features is performed using convolutional operations followed by high-frequency amplifier and low-frequency damping operations. The formula for this stage, including weight matrices, is as follows:
(4) $\mathbf{B}_{Initial}^{l+1}=\mathbf{W}_{B}^{l+1}\odot\widetilde{\mathbf{B}}^{l+1}+\mathbf{W}_{A}^{l}\odot\widetilde{\mathbf{A}}^{l}.$
The $\mathbf{W}_{A}^{l}$ and $\mathbf{W}_{B}^{l+1}$ are the weight matrices associated with the feature $\widetilde{\mathbf{A}}^{l}$ and $\widetilde{\mathbf{B}}^{l+1}$.
Stage 2: Resampling
In Stage 2, the Feature Resampling Module fine-tunes the feature map by adjusting the spatial layout. Features are resampled to achieve better alignment across different layers, ensuring that high-level features are appropriately adjusted to match the finer details of the lower-level features. The formula for this stage is:
(5) $\mathbf{B}_{Intermediate}^{l+1}=Re_{(u,v)}\left(\mathbf{B}_{Initial}^{l+1},\widetilde{\mathbf{B}}^{l+1}\right),$
where $Re_{(u,v)}$ is the resampling operation with displacement $(u,v)$, used to re-align the features from the previous stage, $\mathbf{B}_{Initial}^{l+1}$ is the refined feature from the stage 1, and $\widetilde{\mathbf{B}}^{l+1}$ represents the feature map from the amplifier. Stage 2 refines the spatial alignment of the features, ensuring that they are better aligned for the final fusion stage.
Stage 3: Refined Fusion
In Stage 3, the final fusion of the refined features is performed. This stage combines the results from the previous stages and applies further processing to ensure consistency and clarity of the final output. The formula for this stage is:
(6) $\mathbf{B}^{l}=\mathbf{W}_{B}^{l+1}\odot\mathbf{B}_{Intermediate}^{l+1}+\mathbf{W}_{A}^{l}\odot\widetilde{\mathbf{A}}^{l}.$
Similarity, $\mathbf{W}_{A}^{l}$ and $\mathbf{W}_{B}^{l+1}$ are the weight matrices associated with the feature $\widetilde{\mathbf{A}}^{l}$ and $\mathbf{B}_{Intermediate}^{l+1}$. These weight matrices can be dynamically adjusted based on frequency correlations to ensure better fusion performance. After stage 3, the output leads to the fully refined feature map that can be used as the input to the PHFFNet module of the neck.
3.4. Progressive Hierarchical Feature Fusion Network (PHFFNet)
After introducing FAFCE component, we also propose a novel paradigm PHFFNet for hierarchical feature learning, designed to enhance the object detection model’s ability to learn and align features across different levels. PHFFNet progressively merges features from low to high levels, addressing the significant semantic disparity between non-adjacent layers, especially between the bottom and top features. This gap can weaken feature fusion in autonomous driving contexts, where precise object localization is crucial. By progressively narrowing this semantic gap, the PHFFNet architecture improves feature alignment, optimizing performance for detection tasks in complex driving environments.
Mathematical Representation. We represent the progressive feature fusion process in matrix form, where each feature $C_{n}$ and fused feature $F_{n}$ are expressed as matrices or vectors.
1. Initial Feature Fusion
In the first step, we fuse the low-level features $C_{2}$ and $C_{3}$:
(7) $F_{23}=W_{2,3}\cdot\begin{bmatrix}C_{2}\\ C_{3}\end{bmatrix}.$
Here, $W_{2,3}$ is the weight matrix used for the fusion of low-level features $C_{2}$ and $C_{3}$.
2. Further Fusion
Next, we fuse $F_{23}$ with $C_{4}$, resulting in the new feature $F_{234}$:
(8) $F_{234}=W_{2,3,4}\cdot\begin{bmatrix}F_{23}\\ C_{4}\end{bmatrix},$
where $W_{2,3,4}$ is the weight matrix that facilitates the fusion of $F_{23}$ and $C_{4}$.
3. Final Feature Fusion
Finally, we fuse $F_{234}$ with $C_{5}$, obtaining the final feature $F_{2345}$:
(9) $F_{2345}=W_{2,3,4,5}\cdot\begin{bmatrix}F_{234}\\ C_{5}\end{bmatrix}.$
Here, $W_{2,3,4,5}$ is the weight matrix that facilitates the fusion of $F_{234}$ and $C_{5}$.
We introduce the CASF mechanism within the Hierarchical Fusion Block to assign dynamic spatial weights to features at different levels during multi-level feature fusion. This approach strengthens the capacity of key layers to extract critical information while mitigating the impact of conflicting or irrelevant data from diverse objects and spatial regions, crucial for accurate detection in autonomous driving scenarios.
Let $x_{n}^{\rightarrow l,ij}$ represents the feature vector at spatial position $(i,j)$ from level $n$ to level $l$, and let $y_{l}^{ij}$ denotes the resulting fused feature vector at position $(i,j)$ and level $l$. We integrate feature vectors from 3 distinct Hierarchical Fusion Block as illustrated in Fig. 2. The fusion is performed by summing over these levels, with $\lambda_{n}^{ij}$ representing the dynamic spatial weight for the feature vector at level $n$ and spatial position $(i,j)$. These weights dynamically adjust based on the significance of the features at each spatial location, allowing the fusion process to emphasize relevant features and suppress less important ones.
The fusion is mathematically expressed as:
(10) $y_{l}^{ij}=\sum_{n=1}^{3}\lambda_{n}^{ij}\cdot x_{n}^{\rightarrow l,ij},$
where $\lambda_{n}^{ij}$ are the spatial fusion weights. The constraint